### v2

# On $\operatorname{CAT}(\kappa)$ surfaces

Abstract.
We study the properties of $\operatorname{CAT}(\kappa)$ surfaces: length metric spaces homeomorphic to a surface having curvature bounded above in the sense of satisfying the $\operatorname{CAT}(\kappa)$ condition locally. The main facts about $\operatorname{CAT}(\kappa)$ surfaces seem to be largely a part of mathematical folklore, and this paper is intended to rectify the situation. We provide three distinct proofs that $\operatorname{CAT}(\kappa)$ surfaces have bounded (integral) curvature. This fact allows one to apply the established theory of surfaces of bounded curvature to derive further properties of $\operatorname{CAT}(\kappa)$ surfaces. We also show that $\operatorname{CAT}(\kappa)$ surfaces can be approximated by smooth Riemannian surfaces of Gaussian curvature at most $\kappa$. We do this by giving explicit formulas for smoothing the vertices of model polyhedral surfaces.
Key words and phrases:
Alexandrov geometry, surfaces of bounded curvature, triangulation
2020 Mathematics Subject Classification:
53C45
1. Introduction
The field of Alexandrov geometry concerns two main classes of metric spaces: spaces of curvature bounded below and spaces of curvature bounded above. The latter is known as the class of $\operatorname{CAT}(\kappa)$ spaces, where $\kappa\in\mathbb{R}$ is some real parameter; these can be defined as spaces in which every sufficiently small geodesic triangle is thinner than the triangle of matching edge lengths in the model space of constant curvature $\kappa$. Alexandrov geometry provides a synthetic or coordinate-free approach to the intrinsic geometry of manifolds or more complicated spaces, generalizing ideas from classical Riemannian geometry to potentially non-smooth spaces. $\operatorname{CAT}(\kappa)$ spaces were first studied systematically by Alexandrov [3] in the 1950s, building on his investigations on the intrinsic geometry of convex surfaces; see also earlier work of Busemann [13].
In the two-dimensional setting, there is a third class of spaces that were extensively studied by Alexandrov and the Leningrad school of geometry associated with him, namely surfaces of bounded (integral) curvature. Roughly speaking, this is the largest class of surfaces for which curvature can be meaningfully defined as a signed Radon measure. Its precise definition involves the notion of (angular) excess of a geodesic triangle $T$, denoted by $\delta(T)$. A surface of bounded curvature is one for which, in every compactly contained neighborhood, there is a uniform upper bound on the sum of excesses of an arbitrary finite collection of non-overlapping simple triangles. See Section 2 and Section 3 for complete definitions. There is a rich and well-developed theory of surfaces of bounded curvature developed in the monographs by Alexandrov–Zalgaller [1] and Reshetnyak [38] covering methods such as polyhedral approximation and conformal parametrizations; see also the surveys [19, 41].
We say that a $\operatorname{CAT}(\kappa)$ surface is a metric space that is a topological surface and satisfies the $\operatorname{CAT}(\kappa)$ condition locally. The class of $\operatorname{CAT}(\kappa)$ surfaces includes Riemannian surfaces with Gaussian curvature at most $\kappa$ everywhere, and polyhedral surfaces for which the total angle at each vertex is at least $2\pi$. The purpose of this paper is to give complete and accessible proofs for two fundamental facts concerning $\operatorname{CAT}(\kappa)$ surfaces that, to our knowledge, are not satisfactorily addressed in the literature.
1.1. The $\operatorname{CAT}(\kappa)$ condition and bounded curvature
Our first objective is the following.
Theorem 1.1.
Let $X$ be a $\operatorname{CAT}(\kappa)$ surface. Then $X$ has bounded curvature.
This fact is perhaps not as well known as it could be. For example, the recent papers [8] and [14] prove results directly for $\operatorname{CAT}(\kappa)$ surfaces that are similar to more general results for bounded curvature surfaces, without explicitly recognizing the connection between the two classes of surfaces. 1.1 is trivial if $\kappa\leq 0$, since in this case every triangle has non-positive excess. In the case where $\kappa>0$ its validity is a priori not clear. We note that the analogue of 1.1 for surfaces with curvature bounded below is also true. A proof has been given by Richard in [39].
1.1 was already observed by Alexandrov himself in [3, Section 1.8] in his seminal work on $\operatorname{CAT}(\kappa)$ spaces, justified by an explanation to the effect that one can develop a parallel theory to that of bounded curvature surfaces for surfaces with curvature bounded above by $\kappa$. Alexandrov attributes the result to Zalgaller; however, an elaboration of this approach does not exist in the literature to our knowledge, and in any case seems to be more work than necessary. More recently, a separate proof of 1.1 was sketched by Machigashira–Ohtsuka in [30]. In fact, Lemma 5.3 in [30] states, in more generality, the following basic inequality relating excess and curvature, from which 1.1 follows easily.
Theorem 1.2.
Let $T$ be a triangle in a $\operatorname{CAT}(\kappa)$ surface. Then $\delta(T)\leq\kappa|T|$.
Here, $|T|$ denotes the area of the interior of $T$, which we define as the Hausdorff $2$-measure. According to [30], 1.2 can be established in a similar manner to an analogous inequality for surfaces with curvature bounded below, which was done in earlier work of Machigashira [29, Theorem 2.0]. However, there are various qualitative differences between $\operatorname{CAT}(\kappa)$ spaces and spaces of curvature bounded below, and thus we feel that it is not straightforward to adapt the proof in [29] to the $\operatorname{CAT}(\kappa)$ case.
In light of the state of the literature, Fillastre asked in [17] for complete details of a proof. In this paper, we answer the request by giving three distinct proofs of 1.1, each of which has their own points of interest. These proofs are based on a recent theorem of the third-named author with Creutz [16] on decomposing an arbitrary surface with a length metric into non-overlapping simple convex triangles, which we refer to as a triangulation of the the surface. Note however that the second proof does not require the full strength of the triangulation theorem in [16], but rather the simpler fact that any point has a neighborhood composed of finitely many potentially overlapping triangles (Lemma 5.2 in [16]). The first proof is in our opinion the most comprehensive and intuitive; the main tool we use is a special type of triangulation that can be obtained by inductively subdividing a base triangulation, which we call a vertex-edge triangulation. The second is a more direct geometric proof, intended to establish 1.1 as efficiently as possible using two general facts, namely, Reshetnyak’s majorization theorem and the Besicovitch inequality. This proof is based on proving a non-sharp version of the inequality in 1.2. In the third proof, we verify instead an equivalent definition of being a limit of surfaces of uniformly bounded curvature. This proof is shorter than the others but instead verifies an alternative condition. We attribute the basic idea underlying the third proof to discussions with François Fillastre [18] and a MathOverflow post by Anton Petrunin [36].
Later in the paper, we show how 1.2 itself follows as a consequence of 1.1 using the theory of surfaces of bounded curvature. One technical point relates to the definition of area. As stated above, we take the Hausdorff $2$-measure as our definition. The original work of Alexandrov (cf. [3]) uses a different definition based on approximation by Euclidean polyhedral surfaces, which we call the Alexandrov area. We verify that the two notions of area are equivalent for $\operatorname{CAT}(\kappa)$ surfaces, or more generally for surfaces of bounded curvature without cusp points.
Proposition 1.3.
For any surface of bounded curvature without cusp points, the Alexandrov area coincides with the Hausdorff $2$-measure.
1.2. Smooth approximation
The second objective of this paper concerns approximation by smooth Riemannian surfaces. We prove the following result.
Theorem 1.4.
Let $(X,d)$ be a $\operatorname{CAT}(\kappa)$ surface for some $\kappa\in\mathbb{R}$. Then there exists a sequence of smooth Riemannian metrics $d_{n}$ on $X$ with Gaussian curvature at most $\kappa$ such that the sequence $(X,d_{n})$ converges uniformly to $(X,d)$. Moreover, the metrics $d_{n}$ have the property that $\limsup_{n\to\infty}|A_{n}|=|A|$ for all compact sets $A\subset X$. Here, $A_{n}$ denotes the same set $A$ equipped with the metric $d_{n}$.
While one naturally expects 1.4 to be true, we have not found any complete statement in the literature. From the theory of surfaces of bounded curvature comes the fact that any $\operatorname{CAT}(\kappa)$ surface can be approximated by smooth surfaces with locally uniformly bounded integral curvature, though without any pointwise control on the curvature. See Chapter III of [1] and Section 6.2 of [38]. In [12, Theorem 0.11], Burago–Buyalo show that any $\operatorname{CAT}(\kappa)$ surface can be approximated by piecewise smooth $\operatorname{CAT}(\kappa)$ surfaces (more generally, their result covers any 2-dimensional locally $\operatorname{CAT}(\kappa)$ polyhedron). Recently, the case of 1.4 where $\kappa<0$ has been covered by a lemma of Labeni [24]. Labeni’s argument is based on the approximating cones with vertices of negative curvature in anti-de Sitter space by smooth convex surfaces.
To prove 1.4, the main task is to show that any vertex of a $\operatorname{CAT}(\kappa)$ polyhedral surface can be smoothed while retaining the $\operatorname{CAT}(\kappa)$ condition. We found that this can be accomplished via an explicit and relatively simple formula that interpolates between a metric of constant positive curvature away from the vertex with a metric of constant negative curvature around the vertex.
The analogue of 1.4 for surfaces of curvature bounded below has been addressed in full generality by Itoh–Rouyer–Vîlcu [21]; see also [37, 39] and the discussion in [22, 31] concerning the curvature bounded below case. Our approach also works for smoothing surfaces of curvature bounded below by $\kappa$, with the advantage of being more explicit than the approach in [21]; see 7.2. Thus we provide a concrete and uniform approach to the general problem of smoothing vertices of polyhedra.
1.3. The bigger picture
We now elaborate more on the consequences of 1.1 and place our work within a larger context. While the $\operatorname{CAT}(\kappa)$ condition provides an effective axiomatic condition for studying metric surfaces, there is a variety of increasing general conditions that have also been studied in the literature. Indeed, the $\operatorname{CAT}(\kappa)$ condition is fairly restrictive; roughly speaking, a $\operatorname{CAT}(\kappa)$ surface is one that is the limit of spherical polyhedral surfaces of uniformly bounded area with vertices of negative curvature. We state the following omnibus theorem, which is a compilation of easy or known facts, along with 1.1. For simplicity, we assume that the metric surface $X$ is closed.
Theorem 1.5.
Let $X$ be a closed length surface and $\kappa\in\mathbb{R}$. The following conditions are such that each condition implies the next.
  1. (A)
    $X$ is a smooth Riemannian $2$-manifold with Gaussian curvature at most $\kappa$ everywhere.
  2. (B)
    $X$ is a $\operatorname{CAT}(\kappa)$ surface.
  3. (C)
    $X$ is a surface of bounded curvature without cusp points.
  4. (D)
    $X$ is bi-Lipschitz equivalent to a constant curvature surface of the same topology. In particular, each point of $X$ has a neighborhood bi-Lipschitz equivalent to a Euclidean disk.
  5. (E)
    $X$ is Ahlfors $2$-regular and linearly locally contractible.
  6. (F)
    $X$ satisfies a quadratic isoperimetric inequality.
  7. (G)
    $X$ has locally finite Hausdorff $2$-measure.
Let us define the terms in this theorem that are potentially less familiar or standardized. A metric space $X$ is Ahlfors $2$-regular if there is a constant $C>0$ such that
$\frac{1}{C}\cdot r^{2}\leq|B(x,r)|\leq C\cdot r^{2}$
for all $x\in X$ and $r\in(0,\operatorname{diam}(X))$. Here, $|A|$ denotes the Hausdorff $2$-measure of a set A and $B(x,r)$ is the open ball at $x$ of radius $r>0$. The space $X$ is linearly locally contractible if there exists $C>0$ such that every ball $B(x,r)$ of radius $r\in(0,\operatorname{diam}(X)/C)$ is contractible in $B(x,Cr)$. We say that the space $X$ satisfies a quadratic isoperimetric inequality if every point has a neighborhood $V$ for which there is a constant $C>0$ such that every closed Jordan curve $\Gamma$ in $V$ bounds a topological disk $U\subset V$ satisfying $|U|\leq C\ell(\Gamma)^{2}$. Here $\ell(\Gamma)$ is the length of $\Gamma$.
The implication (A) $\Longrightarrow$ (B) is classical; see the appendix to Chapter II.1 of [9]. The implication (B) $\Longrightarrow$ (C) is a restatement of 1.1 together with the simple observation that a $\operatorname{CAT}(\kappa)$ surface does not have any so-called cusp points, i.e., points having total curvature $2\pi$. Next, by a theorem of Reshetnyak [38, Theorem 9.10], the metric tangent space of every non-cusp point of a surface of bounded curvature is a Euclidean cone over a circle. That is, for each point $x$ on the surface and all $\varepsilon>0$ one can find $\delta>0$ such that the ball $B(x,r)$ is $(1+\varepsilon)$-bi-Lipschitz equivalent to a neighborhood of the above Euclidean cone. See Lemma 6 in [11] for a proof of Reshetnyak’s theorem. See also Bonk–Lang [6] for a strong quantitative version of this property. The implication (C) $\Longrightarrow$ (D) is shown by Burago in [11] using this local bi-Lipschitz equivalence. The implications (D) $\Longrightarrow$ (E) and (F) $\Longrightarrow$ (G) are elementary to verify from the definitions, while the implication (E) $\Longrightarrow$ (F) is due to Lytchak–Wenger [28, Corollary 5.5].
The final condition (G), namely, having locally finite Hausdorff $2$-measure, represents the most general non-fractal metric surface. We briefly remark that fractal surfaces are a separate topic of current interest; these are generally more difficult to study and require separate methods. See for example the monograph of Bonk–Meyer [7]. Condition (G) is very general; nevertheless, the basic idea of the recent papers [16, 34, 35] is that a geometric study of surfaces analogous to that of surface of bounded curvature can be carried out using only the assumption of locally finite Hausdorff $2$-measure. In fact, it turns out that the requirement of a length metric is not so essential and can be omitted.
Regarding condition (F), we mention a beautiful theorem of Lytchak–Wenger [27] that upper curvature bounds can be characterized in terms of isoperimetric inequalities. Namely, a proper length metric space is $\operatorname{CAT}(0)$ if and only if it satisfies the quadratic isoperimetric inequality (defined in a more general way) with the Euclidean constant $1/(4\pi)$. More generally, a proper length metric space is $\operatorname{CAT}(\kappa)$ if and only if its isoperimetric profile is no worse than that of the sphere of constant curvature $\kappa$.
One aspect of the theory concerns the topic of uniformization. The classical uniformization theorem states that any smooth Riemannian $2$-manifold is conformally equivalent to a surface of constant curvature. A major theorem of Reshetnyak concerns conformal parametrizations of bounded curvature surfaces: for any surface of bounded curvature, its metric can be represented locally by a length element of the form $ds=e^{\lambda}ds_{h}$, where $\lambda$ is the difference of two subharmonic functions and $ds_{h}$ is the length element of a background smooth Riemannian metric. The global version of this theorem is due to Huber [20]. There are similar uniformization theorems for the more general classes of surfaces in 1.5, including the Bonk–Kleiner theorem for Ahlfors $2$-regular, linearly locally contractible spheres [5]. A very general uniformization theorem applicable to any surface of locally finite Hausdorff $2$-measure, yielding a so-called weakly quasiconformal parametrization from a constant curvature surface, has been proved in [34, 35].
We refer the reader to the following references for various non-trivial examples of surfaces illustrating the difference between the classes of surfaces in 1.5: Laakso [23] for a surface satisfying (E) but not (D); Theorem 1.1 in [15] for a surface satisfying (F) but not (E); Theorem 1.2 in [40] for a surface satisfying (G) but not (F). In addition, we observe that $\mathbb{R}^{2}$ equipped with any normed metric not arising from an inner product satisfies (D) but not (C). Moreover, any polyhedral surface with cone points of positive curvature satisfies (C) but not (B).
Finally, we remark that there have been several more advanced investigations of finite-dimensional $\operatorname{CAT}(\kappa)$ spaces in recent years; see especially [25, 26]. Two-dimensional metric spaces (not necessarily surfaces) satisfying the $\operatorname{CAT}(\kappa)$ condition have been studied in [32] and [33].
1.4. Outline of the paper
The paper is organized as follows. We collect preliminaries related to metric geometry in Section 2 and related to $\operatorname{CAT}(\kappa)$ spaces in Section 3. The material on $\operatorname{CAT}(\kappa)$ spaces is probably well-known to experts but includes some propositions that we could not find in the literature. Section 4 describes our main tool of vertex-edge triangulations in preparation for the first proof of 1.1. The three proofs of 1.1 are given in Section 5. The other results are proved in the following sections. First, we verify 1.3 on definitions of area and 1.2 relating the excess of a triangle to its area in Section 6. Finally, we prove 1.4 on smooth approximation in Section 7.
Acknowledgments
The research for this project was carried out as part of Stony Brook University’s Research Experience for Undergraduates (REU) program during the summer of 2022. The authors thank the Stony Brook University Mathematics Department for their support. The authors also thank François Fillastre and Alexander Lytchak for their encouragement in completing this project and for feedback on an early draft of this paper. We also especially thank François Fillastre for bringing to our attention several references in the literature.
2. Background
2.1. Basics of metric geometry
We recall the basic definitions related to metric spaces. See standard references such as [9] and [10] for more detail. Let $X$ be a set. A metric on $X$ is a function $d\colon X\times X\to\mathbb{R}$ such that for all $x,y,z\in X$
  1. (1)
    $d(x,y)=0$ if and only if $x=y$ (positive definite),
  2. (2)
    $d(x,y)=d(y,x)$ (reflexive),
  3. (3)
    $d(x,y)\leq d(x,z)+d(z,y)$ (triangle inequality).
The pair $(X,d)$ is called a metric space. Most often, we refer to $X$ itself as a metric space with the convention that $d$ denotes the corresponding metric. We refer to $d(x,y)$ as the distance between $x$ and $y$. Given $x\in X$ and $r>0$, we let $B(x,r)$ denote the open metric ball centered at $x$ of radius $r$. Note that this notation does not specify the metric space, though no confusion should result.
A curve in $X$ is a continuous function from an interval to $X$. The length of a curve $\Gamma\colon[a,b]\to X$ is
$\ell(\Gamma)=\mathrm{sup}\sum_{k=1}^{m}d(\Gamma(t_{k-1}),\Gamma(t_{k})),$
where the supremum is taken over all finite subdivisions $a=t_{0}<t_{1}<\cdots<t_{m}=b$. The curve $\Gamma$ is a geodesic if $\ell(\Gamma)=d(\Gamma(a),\Gamma(b))$. The notation $[xy]$ is used to denote a choice of geodesic from $x$ to $y$, provided that such a geodesic exists. Note that in general such a geodesic need not be unique. In a slight abuse of notation, we also identify a geodesic $[xy]$ with its image.
The metric space $X$ is a length space if
$d(x,y)=\inf_{\Gamma}\ell(\Gamma)$
for all $x,y\in X$, where the infimum is taken over all curves $\Gamma$ connecting $x$ and $y$. The space $X$ is a geodesic space if every pair of points $x,y$ can be joined by a geodesic $[xy]$. Observe in this case that $d(x,y)=\ell([xy])$.
Let $(X,d)$, $(Y,d^{\prime})$ be metric spaces and $L>0$. A map $f\colon X\to Y$ is $L$-Lipschitz if
$d^{\prime}(f(p),f(q))\leq L\cdot d(p,q)$
for all $p,q\in X$. The map $f$ is $L$-bi-Lipschitz if
$\frac{1}{L}d(p,q)\leq d^{\prime}(f(p),f(q))\leq Ld(p,q)$
for all $p,q\in X$.
The diameter of a set $A\subset X$ is
$\operatorname{diam}A=\sup_{x,y\in A}d(x,y).$
The $2$-dimensional Hausdorff $\delta$-content of $A\subset X$ is
$\mathcal{H}_{\delta}^{2}(A)=\mathrm{inf}\left\{\sum_{j=1}^{\infty}c_{2}\cdot\mathrm{diam}(A_{j})^{2}\right\},$
where the infimum is taken over all countable collections of subsets $A_{1},A_{2},\ldots$ of diameter at most $\delta$ such that $A\subset\bigcup_{j=1}^{\infty}A_{j}$. The Hausdorff $2$-measure of $A$ is
$\mathcal{H}^{2}(A)=\lim_{\delta\to 0}\mathcal{H}_{\delta}^{2}(A).$
Here, $c_{2}=\pi/4$ is a normalization constant so that the Hausdorff $2$-measure on the plane coincides with $2$-dimensional Lebesgue measure. Accordingly, we denote the Hausdorff $2$-measure of a set $A$ by $|A|$ rather than $\mathcal{H}^{2}(A)$ throughout this paper. The term “area” in this paper refers specifically to the Hausdorff $2$-measure.
A useful fact about Hausdorff measure is the Besicovitch inequality. We state a two-dimensional version for metric spaces. See Theorem 2.1 in [34] for a proof.
Theorem 2.1 (Besicovitch inequality).
Let $X$ be a metric space homeomorphic to a closed disk with topological boundary $\partial X$. Suppose there is a closed Jordan domain $\Omega\subset\mathbb{R}^{2}$ and a $L$-Lipschitz bijective map from $\partial X$ to $\partial\Omega$. Then
$|X|\geq\frac{\pi|\Omega|}{4L^{2}}.$
2.2. Polygons and triangulations
We start with a general definition of polygon.
Definition 2.2.
Let $X$ be a length metric space and $n\geq 2$ an integer. A (geodesic) $n$-gon is a figure consisting of $n$ points $v_{1},\cdots,v_{n}$, called vertices, together with a collection of $n$ geodesics joining them cyclically, called edges. An $n$-gon is also called a polygon. A polygon is simple if it forms a simple closed curve. A $3$-gon with vertices $x,y,z$ is called a triangle and denoted by $\triangle xyz$. A $4$-gon is called a quadrilateral. The perimeter of a polygon is the sum of the lengths of its edges.
The previous definition is valid for any metric space, although our focus in this paper is two-dimensional metric spaces. A metric surface is a metric space homeomorphic to a topological surface, or $2$-manifold. Consistent with the main references such as [1], we assume that metric surfaces do not have boundary, although in principle the results of this paper should remain true for surfaces with boundary. We use the terms length surface and geodesic surface with the obvious meanings.
We say that a simple polygon $P$ is a Jordan polygon if it is contained in a neighborhood $U$ homeomorphic to an open disk and the diameter of $P$ is at most $d(P,\partial U)$. The last requirement guarantees that the length metric restricted to $P$ is the same in $U$ as in the original space. By the Jordan curve theorem, any simple polygon $P$ is the boundary of a closed Jordan domain contained in the ambient neighborhood $U$. The interior of this domain is called the interior of the polygon, denoted by $P^{\circ}$. Note that the previous fact depends on our assumption of an ambient disk neighborhood; otherwise, a simple polygon may not bound a unique disk, or any disk at all. For example, a great circle on the $2$-sphere can be viewed as a triangle, and either hemisphere may be designated as the interior.
Convention.
When we refer to a “Jordan polygon” $P$, as opposed to simply a “polygon”, we include the interior points of $P$. In turn, we use $\partial P$ to denote the set of edge points of $P$.
Two Jordan polygons are non-overlapping if they have disjoint interiors. A Jordan polygon $P\subset X$ is convex if any two points in $P$ can be joined by a geodesic in $P$. This definition of convexity is sufficient for this paper, although we note that there are stronger notions of convexity that are essential for the study of general length surfaces. See Section 3 of [16] for more discussion.
We state a simple property of convexity.
Proposition 2.3.
Let $P,S$ be two convex Jordan polygons in a metric surface with the property that any two points in $P\cup S$ can be joined by a unique geodesic. If $P^{\circ}\cap S^{\circ}$ is non-empty, then $P\cap S$ is a convex Jordan polygon.
See Lemma 3.2 in [16] for a closely related proposition. Note that the conclusion may fail without the unique geodesic property.
Our proofs make use of covers of a surface by non-overlapping triangles, which we call triangulations. Our definition is somewhat different than the use of the term in general topology, since we do not require adjacent triangles to intersect along entire edges.
Definition 2.4.
A triangulation of a subset $K\subset X$ is a locally finite collection of mutually non-overlapping convex Jordan triangles $\mathcal{T}=\{T_{i}\}$ such that $\bigcup_{i}T_{i}=K$. If the subset $K$ itself is a polygon, then the vertices of $K$ are called the original vertices of the triangulation $\mathcal{T}$. The mesh of $\mathcal{T}$ is $\sup_{T\in\mathcal{T}}\operatorname{diam}T$.
The union of the set of edges of a triangulation $\mathcal{T}$ form an embedded topological graph on the surface, called the edge graph of $\mathcal{T}$ and denoted by $\mathcal{E}$. The edge graph can be considered as a length space by giving it the length metric induced by the metric on $X$.
A general theorem on the existence of triangulations for arbitrary length surfaces was proved in [16]. We state a version that covers our purposes.
Theorem 2.5.
Let $X$ be a length surface and $\mathcal{U}$ a cover of $X$ by topological open disks. For all $\varepsilon>0$ there exists a triangulation $\mathcal{T}$ of $X$ with mesh at most $\varepsilon$ and that each element $T\in\mathcal{T}$ is contained in a set $U\in\mathcal{U}$ satisfying $d(T,\partial U)\geq\varepsilon$. Moreover, if $\mathcal{E}$ is the edge graph of $\mathcal{T}$ with the induced length metric, then the inclusion of $\mathcal{E}$ into $X$ is an $\varepsilon$-isometry.
A map $f\colon X\to Y$ between metric spaces $(X,d)$ and $(Y,d^{\prime})$ is an $\varepsilon$-isometry if
$d(x,y)-\varepsilon\leq d^{\prime}(f(x),f(y))\leq d(x,y)+\varepsilon$
for all $x,y\in X$ and the image of $f$ is $\varepsilon$-dense in $Y$ (that is, every point in $y$ is at distance at most $\varepsilon$ from a point in $f(X)$).
The statement concerning the open cover $\mathcal{U}$ is not part of the triangulation theorem as stated in [16], although it follows from Remark 5.5 in [16]. Also, the triangulation theorem in [16] gives the stronger property of boundary convexity rather than just convexity, though this is not needed for our purposes. The final conclusion regarding the edge graph can be found in Proposition 5.2 in [34].
3. Preliminaries on $\operatorname{CAT}(\kappa)$ spaces
3.1. Basic definitions
In this section, we give a brief overview of $\operatorname{CAT}(\kappa)$ surfaces and collect various facts that are needed later. We refer the reader to Bridson–Haefliger [9] for further details.
For each $\kappa\in\mathbb{R}$, we let $M_{\kappa}^{2}$ denote the $2$-dimensional model space of constant curvature $\kappa$. More precisely, these are defined by letting $M_{0}^{2}$ be the Euclidean plane, $M_{1}^{2}$ be the $2$-sphere, and $M_{-1}^{2}$ be the hyperbolic plane. For $\kappa>0$, $M_{\pm\kappa}^{2}$ is the rescaling of $M_{\pm 1}^{2}$ by the factor $\sqrt{\kappa}$. We let $\bar{d}_{\kappa}$ denote the metric on $M_{\kappa}^{2}$ and $D_{\kappa}$ denote the diameter of $M_{\kappa}^{2}$; that is, $D_{\kappa}=\pi/\sqrt{\kappa}$ if $\kappa>0$ and $D_{\kappa}=\infty$ otherwise. It is often convenient to think of each model space as a conformal deformation of the Euclidean metric, that is, a smooth Riemannian metric obtained by rescaling pointwise the Euclidean metric on a domain in $\mathbb{R}^{2}$. Let $ds_{\kappa}$ denote the corresponding length element for $M_{\kappa}^{2}$; then $ds_{\kappa}$ is given by the formula
$ds_{\kappa}^{2}=\frac{4(dx^{2}+dy^{2})}{(1+\kappa(x^{2}+y^{2}))^{2}}$
defined for all $(x,y)\in\mathbb{R}^{2}$, with the additional requirement that $\sqrt{x^{2}+y^{2}}<1/|\kappa|$ if $\kappa<0$. If $\kappa=0$, this formula gives a rescaling of the usual Euclidean metric, but the resulting metric space is of course isometric to the Euclidean plane. If $\kappa>0$, then we must add in the point $\infty$ to obtain the complete $2$-sphere. One immediate consequence is that any two model spaces are locally bi-Lipschitz equivalent. More precisely, let $\kappa_{1},\kappa_{2}\in\mathbb{R}$, $x_{1}\in M_{\kappa_{1}}^{2}$ and $x_{2}\in M_{\kappa_{2}}^{2}$, and $0<r<\min\{D_{\kappa_{1}},D_{\kappa_{2}}\}$. Then the balls $B(x_{1},r)$ and $B(x_{2},r)$ are $L$-bi-Lipschitz equivalent, where $L\to 1$ uniformly (after fixing $\kappa_{1},\kappa_{2}$) as $r\to 0$. In particular, for all $x_{1}\in M_{0}^{2}$ and $x_{2}\in M_{1}^{2}$, the ball $B(x_{1},2)\subset M_{0}^{2}$ is $2$-bi-Lipschitz equivalent to the ball $B(x_{2},\pi/2)\subset M_{1}^{2}$.
Next, let $X$ be a metric space and fix $\kappa\in\mathbb{R}$. Given a triangle $T=\triangle pqr\subset X$ with perimeter at most $2D_{\kappa}$, a comparison triangle in $M^{2}_{\kappa}$ for $T$ is a triangle in $M^{2}_{\kappa}$ with edges of matching lengths. We denote the comparison triangle of $T$ by $\bar{T}_{\kappa}$. More precisely, let $\bar{p},\bar{q},\bar{r}\in M_{\kappa}^{2}$ be points satisfying $d(p,q)=\bar{d}_{\kappa}(\bar{p},\bar{q})$, $d(p,r)=\bar{d}_{\kappa}(\bar{p},\bar{r})$ and $d(q,r)=\bar{d}_{\kappa}(\bar{q},\bar{r})$. Then we define $\bar{T}_{\kappa}=\triangle\bar{p}\bar{q}\bar{r}$. This triangle is unique up to an isometry of $M^{2}_{\kappa}$.
Given a point $x\in[pq]$, we let $\bar{x}$ denote the point on $[\bar{p}\bar{q}]$ such that $d(p,x)=\bar{d}_{\kappa}(\bar{p},\bar{x})$ and $d(x,q)=\bar{d}_{\kappa}(\bar{x},\bar{q})$. A triangle $T=\triangle pqr\subset X$ with perimeter at most $2D_{\kappa}$ is said to satisfy the $\operatorname{CAT}(\kappa$) inequality if
$d(x,y)\leq\bar{d}_{\kappa}(\bar{x},\bar{y})$
for all $x,y$ in $\partial T$. A length metric space $X$ is a $\operatorname{CAT}(\kappa)$ space if any two points in $X$ at distance less than $D_{\kappa}$ are joined by a geodesic and every triangle $T\subset X$ with perimeter less than $2D_{\kappa}$ satisfies the CAT($\kappa$) inequality. We say that a length surface $X$ is a locally $\operatorname{CAT}(\kappa)$ surface if every point $x\in X$ has a neighborhood $U$ that is a $\operatorname{CAT}(\kappa)$ space as previously defined. For conciseness, we typically omit the word locally and refer to such a space as a $\operatorname{CAT}(\kappa)$ surface.
3.2. Angles
Let $X$ be an arbitrary metric space. Given a triangle $T=\triangle pqr\subset X$, the model angle at $p$, denoted ¯$\bar{\angle}_{p}^{\kappa}(q,r)$, is the angle between the geodesics $[\bar{p}\bar{q}]$ and $[\bar{p}\bar{r}]$ in the comparison triangle $\triangle\bar{p}\bar{q}\bar{r}$ in $M^{2}_{\kappa}$. This is defined provided that the perimeter of $T$ is at most $2D_{\kappa}$. In a $\operatorname{CAT}(\kappa)$ space, the model angle satisfies the triangle inequality: ¯$\bar{\angle}_{p}^{\kappa}(q,r)\leq\bar{\angle}_{p}^{\kappa}(q,s)+\bar{\angle}_{p}^{\kappa}(s,r)$ for all $p,q,r,s$, provided that all these angles are defined.
For any metric space $X$, the (Alexandrov) upper angle between geodesics $[pq]$ and $[pr]$ is
(3.1) $\angle_{p}(q,r)=\limsup_{q^{\prime},r^{\prime}\to p}\bar{\angle}_{p}^{0}(q^{\prime},r^{\prime})$
[…]
Let $T$ be a triangle with angles $\alpha,\beta,\gamma$. The angle excess or excess of $T$ is $\delta(T)=\alpha+\beta+\gamma-\pi$. If $\mathcal{T}$ is a finite family of geodesic triangles, the excess of $\mathcal{T}$ is $\delta(\mathcal{T})=\sum_{i}\delta(T_{i})$. In connection with excess, we recall the classical Girard’s theorem, which states that the area of a triangle in the $2$-sphere $M_{1}^{2}$ is precisely equal to its excess. More generally, if $T$ is a triangle in the model space $M_{\kappa}^{2}$ for $\kappa>0$, then $|T|=\delta(T)/\kappa$.
3.3. Spherical law of cosines
In this section, we restrict to the case where $\kappa>0$. We recall the spherical law of cosines. See for example Section 1.A of [2]. Consider a triangle with side lengths $a,b,c$ and corresponding opposite angles $\alpha,\beta,\gamma$ in the model space $M_{\kappa}^{2}$. Then
$\cos\gamma=\frac{\cos(c\sqrt{\kappa})-\cos(a\sqrt{\kappa})\cos(b\sqrt{\kappa})}{\sin(a\sqrt{\kappa})\sin(b\sqrt{\kappa})}.$
By expanding this equation with power series, we obtain the following inequality for fixed $\kappa>0$: for all $\varepsilon>0$, there exists $\delta>0$ such that if $a,b<\delta$ then
(3.2) $(1-\varepsilon)(a^{2}+b^{2}-2ab\cos\gamma)\leq c^{2}\leq(1+\varepsilon)(a^{2}+b^{2}-2ab\cos\gamma).$
As a particular consequence, if $\gamma\geq\theta$ for some fixed value $\theta>0$, then there is a corresponding constant $C_{\theta}>0$ such that
(3.3) $C_{\theta}\min\{a^{2},b^{2}\}\leq(1-\varepsilon)(a^{2}+b^{2}-2ab\cos\gamma)\leq c^{2}$
whenever $a,b<\delta$.
The following lemma forms the basis of the second proof of 1.1.
Lemma 3.1.
Fix $\kappa>0$. For all $\theta>0$, there exists $L\geq 1,\delta>0$ such that the following holds for every triangle $T$ in a $\operatorname{CAT}(\kappa)$ space with each angle at least $\theta$ and having perimeter at most $\delta$.
  1. (1)
    The ratio of the length of any two edges of $T$ is bounded by $L$.
  2. (2)
    The canonical map from $T$ to its Euclidean model triangle $\bar{T}_{0}$ is $L$-bi-Lipschitz.
  3. (3)
    If $T$ is a Jordan triangle, then
    $\delta(T)\leq 4|\bar{T}_{0}|\leq 16L^{2}|T|.$
The canonical map in (2) is the map that sends each edge of $T$ to the corresponding edge of $\bar{T}_{0}$ preserving arc length, that is, each point $x\in T$ to the point $\bar{x}\in\bar{T}_{0}$.
Proof.
By rescaling, we may assume that $\kappa=1$. Choose $\delta>0$ small enough so that (3.2) holds for $\varepsilon=1/2$ for any triangle in $M_{\kappa}^{2}$ with perimeter at most $\delta$. Assume further that $\delta\leq\pi/2$. Consider a triangle $T$ with vertices $p,q,r$, with each angle at least $\theta$ and perimeter at most $\delta$. Statement (1) follows immediately from (3.3).
We now verify statement (2). Let $\bar{T}_{1}$ be the spherical model triangle corresponding to $T$ and let $\bar{T}_{0}$ be the Euclidean model triangle. Let $x,y\in T$. If $x,y$ belong to the same side, then $d(x,y)=\bar{d}_{1}(\bar{x},\bar{y})$ and there is nothing more to prove. Otherwise, by symmetry, we may assume that the sides containing $x$ and $y$, respectively, meet at a common vertex $p$. Let $\theta_{0}$ denote the angle at the vertex $\bar{p}$ in $\bar{T}_{1}$. Let $\bar{T}=\triangle\bar{p}^{\prime}\bar{x}^{\prime}\bar{y}^{\prime}$ denote the spherical model triangle corresponding to the triangle $\triangle pxy$, and $\bar{\theta}$ the angle at $\bar{p}^{\prime}$. The $\operatorname{CAT}(\kappa)$ condition implies
$d(x,y)=\bar{d}_{1}(\bar{x}^{\prime},\bar{y}^{\prime})\leq\bar{d}_{1}(\bar{x},\bar{y})$
and moreover that $\theta_{0}\geq\bar{\theta}\geq\theta$. For the following inequalities, we use the auxiliary function
$L(\rho)=d(p,x)^{2}+d(p,y)^{2}-2d(p,x)d(p,y)\cos(\rho).$
Applying (3.2) with the value $\varepsilon=1/2$ gives
$(1/2)L(\bar{\theta})\leq\bar{d}_{1}(\bar{x}^{\prime},\bar{y}^{\prime})^{2}=d(x,y)^{2}.$
Since $\theta\leq\bar{\theta}$, the left-hand side of the above inequality is bounded below by $(1/2)L(\theta)$, which in turn has a lower bound of the form $C_{\theta}^{\prime}(d(p,x)^{2}+d(p,y)^{2})$, $C_{\theta}^{\prime}$ a constant depending only on $\theta$. On the other hand, we have the upper bound $L(\theta_{0})\leq d(p,x)^{2}+d(p,y)^{2}$. By (3.2), we have $L(\theta_{0})\geq(2/3)\bar{d}_{1}(\bar{x},\bar{y})^{2}$. Combining these inequalities gives
$\frac{L(\bar{\theta})}{2}=\frac{L(\bar{\theta})}{2L(\theta_{0})}L(\theta_{0})\geq\frac{C_{\theta}^{\prime}}{2}\cdot\frac{2}{3}\bar{d}_{1}(\bar{x},\bar{y})^{2}.$
We conclude that $\bar{d}_{1}(\bar{x},\bar{y})\leq\sqrt{3/C_{\theta}^{\prime}}\cdot d(x,y)$, hence that the canonical map from $T$ to $\bar{T}_{1}$ is bi-Lipschitz with constant only depending on $\theta$. Observe that, since the perimeter of $T$ is at most $\pi/2$, the triangles $\bar{T}_{0}$ and $\bar{T}_{1}$ are $2$-bi-Lipschitz equivalent. Then (2) follows immediately.
Finally, we show that (3) holds as a consequence of the Besicovitch inequality (2.1). Recall that the triangles $\bar{T}_{0}$ and $\bar{T}_{1}$ are $2$-bi-Lipschitz equivalent. Applying the Besicovitch inequality and Girard’s theorem, we see that
$|T|\geq\frac{|\bar{T}_{0}|}{4L^{2}}\geq\frac{|\bar{T}_{1}|}{16L^{2}}=\frac{\delta(\bar{T}_{1})}{16L^{2}}\geq\frac{\delta(T)}{16L^{2}}.$
This verifies (3). ∎
A similar statement, with more intricate proof, holds for surfaces of bounded curvature. See Lemma 4 in [4].
3.4. Facts about $\operatorname{CAT}(\kappa)$ spaces
We record a few foundational facts about $\operatorname{CAT}(\kappa)$ spaces. The first of these is Reshetnyak’s majorization theorem; see Section 9.L of [2].
Let $D\subset M_{\kappa}^{2}$ be a closed Jordan domain whose boundary has finite length, and let $X$ be a metric space. A map $f\colon D\to X$ is majorizing if it is $1$-Lipschitz and its restriction to the boundary $\partial D$ is length-preserving. If $\Gamma$ is a closed curve in $X$, we say that $D$ majorizes $\Gamma$ if there is a majorizing map $f\colon D\to X$ such that the restriction $f|_{\partial D}$ traces out $\Gamma$.
Theorem 3.2 (Reshetnyak’s majorization theorem).
For any closed curve $\Gamma$ in a $\operatorname{CAT}(\kappa)$ space of length at most $2D_{\kappa}$, there exists a convex region $D$ in $M^{2}_{\kappa}$, and an associated map $f$ such that $D$ majorizes $\Gamma$ under $f$.
If $\Gamma$ itself is the boundary of a simple triangle $T$, then it is easy to see that $D$ is the model triangle $\bar{T}_{\kappa}$.
A consequence of Reshetnyak’s majorization theorem is that any $\operatorname{CAT}(\kappa)$ surface has locally finite Hausdorff $2$-measure. Consider a triangulation $\mathcal{T}$ of the surface as given by 2.5, where $\varepsilon$ is chosen to satisfy $\varepsilon<D_{\kappa}/2$ and each triangle is contained in a $\operatorname{CAT}(\kappa)$ neighborhood. For each triangle $T\in\mathcal{T}$, Reshetnyak’s theorem gives a $1$-Lipschitz map $\bar{T}_{1}\to T$. Since $\bar{T}_{1}$ has finite area and $\mathcal{T}$ is locally finite, we see that the Hausdorff $2$-measure on $X$ is locally finite. Note however that we do not need the full strength of the triangulation theorem for this conclusion, but rather just the much easier property that any point lies in a polygonal neighborhood (Lemma 5.1 in [16]).
The next fact about $\operatorname{CAT}(\kappa)$ spaces concerns uniqueness of geodesics. If two points $x,y\in X$ satisfying $d(x,y)<D_{\kappa}$ are joined by a geodesic, then that geodesic is unique. The existence of a geodesic between sufficiently close points is a consequence of the Hopf–Rinow theorem; see Proposition 3.7 in [9]. Suppose that $x,y$ belong to a neighborhood $U\subset X$ homeomorphic to a disk. If $d(x,y)\leq d(x,\partial U)$, then there is a geodesic in $X$ from $x$ to $y$. Moreover, the ball $B(x,r)$ is convex and contractible for all $r<\min\{d(x,\partial U)/2,D_{\kappa}/2\}$.
Lemma 3.3.
Let $X$ be a $\operatorname{CAT}(\kappa)$ space that is also a topological manifold. Then every geodesic can be extended indefinitely in both directions as a local geodesic. It is length-minimizing for all pairs of points at distance less than $D_{\kappa}$.
See Proposition II.5.125.125.125.12 in [9] for a proof.
Corollary 3.4.
Let $T=\triangle pqr$ be a Jordan triangle and $x$ an interior point. Then the geodesic $[px]$ can be extended to intersect the opposite side $[qr]$.
Proof.
By 3.3, $[px]$ can be extended indefinitely beyond $x$. Since the triangle $T$ is compact, this extension must intersect the boundary $\partial T$ at some point. This point must lie on the side $[yz]$ since otherwise the uniqueness of geodesic is violated. ∎
As a consequence of geodesic extendability, we can show that small triangles in $\operatorname{CAT}(\kappa)$ spaces are convex.
Lemma 3.5.
Let $T$ be a Jordan triangle in a $\operatorname{CAT}(\kappa)$ space with perimeter at most $D_{\kappa}/2$. Assume further that $T$ is contained in a neighborhood $U$ homeomorphic to an open disk with $d(T,\partial U)\geq\ell(\partial T)$. Then $T$ is convex.
Proof.
Consider two points $x,y\in T$. Let $\Gamma$ be the geodesic between them, which exists by the second assumption. We must show that the image of $\Gamma$ is contained in $T$. If $x,y$ belong to the same side of $T$, then $\Gamma$ is just a subarc of this side and we are done. Assume then that $x$ belongs to the side $A$ and $y$ belongs to the side $B$, with $C$ the other side of $T$. Extend the side $A$ in both directions until reaching $\partial U$. This divides $U$ into two non-overlapping closed disks $U_{A}^{1}$ and $U_{A}^{2}$, where $T\subset U_{A}^{1}$. Do the same for $B$ to obtain two non-overlapping closed disks $U_{B}^{1}$ and $U_{B}^{2}$. If $\Gamma$ leaves the set $U_{A}^{1}$, then we homotope any subcurves of $\Gamma$ in $(U_{A}^{2})^{\circ}$ onto $A$ to obtain a curve $\Gamma_{1}$ in $U_{A}^{1}$. It is clear that $\ell(\Gamma_{1})\leq\ell(\Gamma)$. Next, if $\Gamma_{1}$ leaves the set $U_{B}^{1}$, we do the same procedure to obtain a curve $\Gamma_{2}$ in $U_{B}^{1}\cap U_{A}^{1}$. This curve necessarily satisfies $\ell(\Gamma_{2})\leq\ell(\Gamma_{1})$. Finally, if the curve $\Gamma_{2}$ leaves $T$, it must be through the side $C$. We homotope any subcurve of $\Gamma_{2}$ in $U\setminus T$ onto $C$ to obtain a curve $\Gamma_{3}$ contained in $T$. It is clear that $\ell(\Gamma_{3})\leq\ell(\Gamma_{2})$. ∎
3.5. Surfaces of bounded curvature
In this section, we give a brief overview of the theory of surfaces of bounded curvature. See the monographs [1] and [38] for a detailed treatment, as well as [41] for a shorter survey. The fundamental definition in the theory as developed by Alexandrov–Zalgaller is the following.
Definition 3.6.
A length surface $X$ has bounded curvature if for each $x\in X$ there is a neighborhood $U$ containing $x$ and a constant $C>0$ such that $\delta(\mathcal{T})\leq C$ for all finite collections $\mathcal{T}$ of mutually non-overlapping Jordan triangles contained in $U$.
The main method of proof used by Alexandrov–Zalgaller is polyhedral approximation. For a flat polyhedral surface $Y$ (i.e., a polyhedral surface where each face is a Euclidean polygon), one can define the total curvature at a vertex to be $2\pi$ minus the total angle at that vertex. One can then define the curvature measure $\omega$ on $Y$ as a signed Radon measure by defining $\omega(A)$ to be the sum of the total curvatures of all vertices in $A$. The measure $\omega$ in turn can be split into a positive part $\omega^{+}$ and negative part $\omega^{-}$, where $\omega^{+}$ and $\omega^{-}$ are non-signed measures satisfying $\omega=\omega^{+}-\omega^{-}$. Suppose that $(X,d)$ is a length surface and $(d_{n})$ is a sequence of polyhedral metrics on $X$ converging uniformly to $d$, with $\omega_{n}$ the corresponding curvature measures. We say that the sequence of metrics $d_{n}$ has locally uniformly bounded curvature if for any compact set $A\subset X$ both $\omega_{n}^{+}(A)$ and $\omega_{n}^{-}(A)$ are uniformly bounded in $n$.
Surfaces of bounded curvature can be defined equivalently in terms of a sequence of polyhedral approximations. Namely, a length surface $(X,d)$ has bounded curvature if and only if there is a sequence of polyhedral metrics $d_{n}$ on $X$ with locally uniformly bounded curvature that converges uniformly to $d$. The same statement is true with “polyhedral metric” replaced by “Riemannian metric”; in this case, the curvature measure is defined by integrating the Gaussian curvature with respect to area. Surfaces of bounded curvature can also be characterized in terms of the existence of conformal parametrizations due to Reshetnyak, which we discussed briefly in Section 1.3.
The characterization as limits of polyhedral surfaces enables one to define area and the curvature measure on a surface of bounded curvature by taking weak limits. We call the notion of area obtained this way the Alexandrov area and denote the area of a set $A$ by $\operatorname{Area}(A)$. More precisely, let $(d_{n})$ be a sequence of polyhedral metrics on $X$ with locally uniformly bounded curvature converging uniformly to a metric $d$, with $\varphi_{n}$ denoting the transition map from $(X,d_{n})$ to $(X,d)$ and $\omega_{n}$ the curvature measure for $d_{n}$. For all $A\subset X$ we define $\operatorname{Area}(A)=\lim_{n\to\infty}|\varphi_{n}^{-1}(A)|$. Likewise, we define $\omega(A)=\lim_{n\to\infty}\omega_{n}(\varphi_{n}^{-1}(A))$ It can be shown that these are independent of the choice of approximating surface and hence well-defined; see Section VIII.3 of [1].
We mention a few more advanced concepts related to surfaces of bounded curvature. First is the notion of sector angle. Let $p\in X$ be a point, let $[pq]$ and $[pr]$ be geodesics intersecting only at $p$, and let $O$ be a closed oriented Jordan neighborhood of $p$ divided into two sectors $U$ and $V$. One of these, say $U$, contains the arc of $\partial O$ running counterclockwise from $[pq]$ to $[pr]$. We define the sector angle of $[pq]$ and $[pr]$ as follows. Let $a_{0},\ldots,a_{n}$ be a sequence of points in $\partial U\cap\partial O$ in cyclic order with $a_{0}\in[pq]$ and $a_{n}\in[pr]$, and $[pa_{1}],\ldots,[pa_{n}]$ be a sequence of geodesics intersecting $\partial U$ only at their endpoints and intersecting each other only at $p$. The sector angle of $[pq]$ and $[pr]$, denoted by ~$\widetilde{\angle}_{p}(q,r)$, is the supremum of the sum $\sum_{i=1}^{n}\angle_{p}(a_{i},a_{i+1})$, where the supremum is taken over all such $n$-tuples $a_{0},\ldots,a_{n}$ and admissible choices of geodesics. The sector angle is independent of the choice of neighborhood $O$.
Next, the total angle at a point p is the supremum of the sum of sector angles taken over all partitions of a neighborhood of $p$ into sectors. We record the observation that if $p$ is the interior point of a geodesic, then the sector angle at $p$ is at least $2\pi$, since the sector corresponding to each side of the geodesic contributes at least $\pi$ to the total angle.
Let $X$ be a surface of bounded curvature and let $D$ be a Jordan polygon. Denote the sector angle at a vertex $v_{i}$ of $\partial D$ on the interior side of $\partial D$ by $\phi_{i}$. Assume further that the interior sector angle of each non-vertex point of $\partial D$ is $\pi$. We define the interior rotation of the curve $\partial D$ by $\tau(\partial D)=\sum_{i}(\pi-\phi_{i})$. The interior rotation can be defined for more general simple closed curves by taking limits of polygons, though we do not need this generality here.
The following version of the Gauss–Bonnet theorem holds for surfaces of bounded curvature. See Theorem VI.5 of [1].
Theorem 3.7 (Gauss–Bonnet).
If $X$ is a surface of bounded curvature and $D\subset X$ is a polygon, then
$\tau(\partial D)+\omega(D^{\circ})=2\pi.$
4. Vertex-Edge Triangulations
In this section, we define vertex-edge triangulations and establish their main properties in preparation for the first proof of 1.1. Throughout this section, $X$ is a $\operatorname{CAT}(\kappa)$ surface.
All polygons we consider for the remainder of the paper are Jordan polygons, so we drop the word “Jordan” and just use the term “polygon”. We recall our convention that interior points are considered part of a (Jordan) polygon.
First, we formally state the relevant definitions.
Definition 4.1.
Let $P\subset X$ be a convex polygon. A vertex-edge partition of $P$ is a cover of $P$ by non-overlapping convex polygons that can be obtained by the following inductive process. First, $\mathcal{T}_{0}=\{P\}$ itself is a vertex-edge partition.
  • Vertex-edge subdivision. Let $\mathcal{T}$ be a vertex-edge partition. Given a polygon $T\in\mathcal{T}$, we subdivide it by connecting one of its vertices to a point on an edge not containing that vertex by a geodesic in $T$ that intersects $\partial T$ only at the endpoints. This produces two non-overlapping convex polygons $T_{1}$ and $T_{2}$ whose union is $T$. We define a new vertex-edge partition $\mathcal{T}^{\prime}$ by replacing $T$ with the two polygons $T_{1},T_{2}$.
If a vertex-edge partition $\mathcal{T}$ of $P$ consists only of triangles, then we call it a vertex-edge triangulation of $P$.
We make the following useful observation: if $P$ is a triangle, then every intermediate polygon appearing in the vertex-edge subdivision is also a triangle.
Definition 4.2.
Suppose that $\mathcal{A}$ is a family of non-overlapping convex polygons contained in a polygon $P$. A vertex-edge refinement of $\mathcal{A}$ with respect to $P$ is a triangulation $\mathcal{T}$ with the following properties:
  1. (1)
    $\mathcal{T}$ is a vertex-edge triangulation of $P$;
  2. (2)
    Each polygon $A\in\mathcal{A}$ is the union of triangles in $\mathcal{T}$.
  3. (3)
    For each $A\in\mathcal{A}$, the set $\mathcal{T}_{A}=\{T\in\mathcal{T}:T\subset A\}$ is itself a vertex-edge triangulation of $A$.
The point of this definition is that $\mathcal{T}$ is simultaneously a vertex-edge triangulation of the original triangle $T$ and each of the polygons in $\mathcal{A}$.
4.1. Existence of vertex-edge refinements
First, we have a lemma about recognizing vertex-edge triangulations.
Lemma 4.3.
Let $P$ be a polygon in a $\operatorname{CAT}(\kappa)$ surface of perimeter less than $2D_{\kappa}$ such that $P$ is the union of non-overlapping convex polygons $P_{1},\ldots,P_{n}$, where each $P_{i}$ is a triangle or quadrilateral. Assume that each polygon $P_{i}$ intersects $P_{i+1}$ along a single common edge, and that the $P_{i}$ do not intersect otherwise except at vertices. Moreover, if $P_{i}$ is a quadrilateral, then we assume that it does not intersect $P_{i-1}$ and $P_{i+1}$ along adjacent edges of $P_{i}$. Finally, we require each original vertex of $P$ to be a vertex of one of the $P_{i}$. We form a triangulation $\mathcal{T}$ of $P$ by dividing each $P_{i}$ that is a quadrilateral into triangles by connecting two non-adjacent vertices with a geodesic. Then $\mathcal{T}$ is a vertex-edge triangulation of $P$.
Proof.
Observe first that the perimeter assumption on $P$ implies that $P$ has at least three original vertices.
The lemma follows by induction on $n$, the number of subpolygons comprising $P$. It is clear that the conclusion holds if $n=1$. Assume next that the conclusion holds whenever $n\leq N-1$ for some integer $N\geq 2$. Let $P$ be a polygon comprised of $n=N$ subpolygons as in the statement of the lemma. In the first case, suppose that $P$ has an original vertex that is a vertex of a polygon $P_{i}$ for some $i\in\{2,\ldots,n-1\}$. This vertex must be contained in a common edge of $P_{i}$ with $P_{i-1}$ or $P_{i+1}$. Subdivide $P$ along this common edge, giving two subpolygons handled by the inductive hypothesis.
Otherwise, either $P_{1}$ or $P_{n}$ is a quadrilateral containing two original vertices of $P$ not in common with $P_{2}$ or $P_{n-1}$, respectively. Say that $P_{1}$ has this property. Let $e$ be the edge of $\mathcal{T}$ containing one of these original vertices and crossing through $P$; then $e$ must connect one original vertex of $P$ and one vertex of the common edge with $P_{2}$. So we subdivide $P$ first along $e$, then along the common edge with $P_{2}$. This divides $P$ into the two subtriangles of $P_{1}$ and the polygon $P^{\prime}=P_{2}\cup\cdots\cup P_{n}$, which is handled by the inductive hypothesis. ∎
The main result of this section is the following.
Theorem 4.4.
If $T$ is a convex triangle of perimeter less than $2D_{\kappa}$ that is a $\operatorname{CAT}(\kappa)$ space, then any finite family $\mathcal{A}$ of non-overlapping convex polygons contained in $T$ has a vertex-edge refinement with respect to $T$.
The proof is somewhat technical, but the geometric idea at its core can be easily read off from Figure 4.1
Proof.
Let $T=\triangle pqr$ be such a triangle and let $\mathcal{A}=\{A_{i}\}_{i=1}^{n}$ be a finite collection of non-overlapping convex polygons in $T$. Enumerate the vertices of the $A_{i}$ as $\{v_{j}\}_{j=1}^{m}$. We give the following procedure to find a vertex-edge refinement. Note that our assumption on $T$ implies that any pair of points in $T$ are connected by a unique geodesic.
  1. (1)
    For each vertex $v_{j}$ not contained in $[pq]$ or $[pr]$, extend the unique geodesic $[pv_{j}]$ to a geodesic that intersects the opposite edge $[qr]$ using 3.3. See Figure 1(b). This yields a collection of distinct geodesics $\{I_{k}\}_{k=1}^{K}$, enumerated in increasing order based upon where the endpoint lies on $[qr]$, oriented from $q$ to $r$. Note that two geodesics $I_{j_{1}},I_{j_{2}}$ cannot share the same endpoint, since otherwise they must coincide by uniqueness of geodesics. Also, set $I_{0}=[pq]$ and $I_{K+1}=[pr]$. For all $k\in\{0,\ldots,K\}$, the geodesics $I_{k},I_{k+1}$ together with a subarc of $[qr]$ bound a triangle $T_{k}$.
  2. (2)
    For each triangle $T_{k}$ apply the following. Let $\mathcal{A}_{k}$ denote the collection of polygons of the form $A_{i}\cap T_{k}$ intersecting the interior of $T_{k}$; denote this polygon by $T_{k}^{i}$. By 2.3, $T_{k}^{i}$ is a convex polygon. Note that each vertex of each $T_{k}^{i}$ is contained in $I_{k}\cup I_{k+1}$. This implies that every edge of some $T_{k}^{i}$ not contained in $I_{k}$ or $I_{k+1}$ must connect $I_{k}$ to $I_{k+1}$. Enumerate such distinct edges as $\{J_{k}^{j}\}_{j=1}^{J_{k}}$ in increasing order from $[qr]$ to the vertex $p$. Also, denote the subarc of $[qr]$ contained in $T_{k}$ by $J_{k}^{0}$. Note that $J_{k}^{j}$ and $J_{k}^{j+1}$ along with the relevant subarcs of $I_{k}$ and $I_{k+1}$ form either a triangle or quadrilateral. We subdivide each triangle $T_{k}$ as follows. For all $j\in\{0,\ldots,J_{k}-1\}$, if $J_{k}^{j}$ and $J_{k}^{j+1}$ do not share an endpoint, then connect an endpoint of $J_{k}^{j}$ to an endpoint of $J_{k}^{j+1}$ by a geodesic, forming two triangles. In this manner we obtain a triangulation $\mathcal{T}$ of $P$. The final product is shown in Figure 1(c).
[…]
Figure 4.1.
It follows from construction that $\mathcal{T}$ is a vertex-edge triangulation of $T$. Indeed, $\mathcal{T}$ is obtained first by drawing geodesics from $p$ to the opposite edge to form the subtriangles $T_{k}$. Then each $T_{k}$ is the union of triangles satisfying the requirements of 4.3. We conclude that $\mathcal{T}$ is vertex-edge.
The final property is that, for each $A\in\mathcal{A}$, the collection $\mathcal{T}_{A}$ can be obtained by vertex-edge subdivision of $A$. Observe that the geodesics $\{I_{k}\}_{k=1}^{K}$ divide $A$ into the non-overlapping union of triangles and quadrilaterals satisfying the requirements of 4.3. We now apply this lemma to conclude that $\mathcal{T}_{A}$ is a vertex-edge triangulation of $A$. ∎
4.2. Subdividing triangles
The main usefulness of vertex-edge triangulations stems from the following lemma, which states that both excess and model area behave monotonically when subdividing triangles.
Lemma 4.5.
Let $T=\Delta pqr$ be a triangle of perimeter at most $2D_{\kappa}$ satisfying the $\operatorname{CAT}(\kappa)$ condition and let $s$ be a point on the side $[qr]$. Let $\mathcal{T}$ be the vertex-edge triangulation consisting of the triangles $T_{1}=\Delta psq$ and $T_{2}=\Delta psr$ formed by connecting $p$ and $s$. then
  1. (1)
    Total excess does not decrease:
    $\delta(T)\leq\delta(T_{1})+\delta(T_{2})$
  2. (2)
    Total model area does not increase:
    $|\bar{T}_{1}|+|\bar{T}_{2}|\leq|\bar{T}|,$
    where $\bar{T},\bar{T}_{1},\bar{T}_{2}$ are the corresponding comparison triangles in the model space $M_{\kappa}^{2}$.
Proof.
Denote the angles of $T$ by $\alpha,\beta,\gamma$ and the angles of each $T_{i}$ by $\alpha_{i},\beta_{i},\gamma_{i}$, where $\alpha$, $\alpha_{1}$, $\alpha_{2}$ are the angles at the shared vertex $p$ and $\gamma_{1}$, $\beta_{2}$ are the angles at the vertex $s$.
We first prove (1). The triangle inequality for angles implies that $\alpha\leq\alpha_{1}+\alpha_{2}$ and $\pi\leq\gamma_{1}+\beta_{2}$. It follows that
$\delta(T)$ $=\alpha+\beta+\gamma-\pi$
$\leq\alpha_{1}+\alpha_{2}+\beta+\gamma+(\gamma_{1}+\beta_{2}-\pi)-\pi$
$=(\alpha_{1}+\gamma_{1}+\beta-\pi)+(\alpha_{2}+\beta_{2}+\gamma-\pi)$
$=\delta(T_{1})+\delta(T_{2}).$
Next, to prove (2), the main task is to show that the comparison triangles $\bar{T}_{1}$ and $\bar{T}_{2}$ can be fitted into $\bar{T}$ such that $\bar{T}_{1}$ and $\bar{T}_{2}$ are non-overlapping. Our argument is based on Alexandrov’s Lemma; see [9, I.2.16]. Consider the triangle $\bar{T}_{1}$ arranged to share the side $[\bar{p}\bar{q}]\subset\bar{T}$ and have overlapping interior. Similarly, the triangle $\bar{T}_{2}$ is arranged to share the side $[\bar{p}\bar{r}]\subset\bar{T}$ and have overlapping interior. Let $\bar{y}$ denote the third vertex of $\bar{T}_{1}$ and let $\bar{z}$ denote the third vertex of $\bar{T}_{2}$. The situation is summarized in Figure 4.2.
[…]
Figure 4.2. Subdividing a $\operatorname{CAT}(\kappa)$ triangle
Let $\bar{\alpha}$, $\bar{\beta}$, $\bar{\gamma}$ denote the angle in $\bar{T}$ corresponding to $\alpha$, $\beta$, $\gamma$, respectively, and likewise for each $\bar{T}_{i}$. The $\operatorname{CAT}(\kappa)$ condition implies that $\gamma_{1}\leq\bar{\gamma}_{1}$ and $\beta_{2}\leq\bar{\beta}_{2}$, and hence $\pi\leq\bar{\gamma}_{1}+\bar{\beta}_{2}$. Then Alexandrov’s Lemma states that $\bar{\alpha}\geq\bar{\alpha}_{1}+\bar{\alpha}_{2}$, $\bar{\beta}_{1}\leq\bar{\beta}$ and $\bar{\gamma}_{2}\leq\bar{\gamma}$.
Since $\bar{\beta}_{1}\leq\bar{\beta}$ and $\alpha_{1}\leq\alpha$, we see that $\bar{y}$ is contained in $\bar{T}$. Likewise, $\bar{z}$ is also contained in $\bar{T}$. This implies that the triangles $\bar{T}_{1}$ and $\bar{T}_{2}$ are entirely contained in $\bar{T}$. Moreover, since $\bar{\alpha}\geq\bar{\alpha}_{1}+\bar{\alpha}_{2}$, the segments $[\bar{p}\bar{r}]$, $[\bar{p}\bar{z}]$, $[\bar{p}\bar{y}]$, $[\bar{p}\bar{q}]$ emanate from $\bar{p}$ counterclockwise in this order, with possibly $[\bar{p}\bar{z}]$, $[\bar{p}\bar{y}]$ coinciding. From this it follows that $\bar{T}_{1}$ and $\bar{T}_{2}$ do not overlap. We conclude that $|\bar{T}_{1}|+|\bar{T}_{2}|\leq|\bar{T}|$. ∎
5. $\operatorname{CAT}(\kappa)$ surfaces have bounded curvature
5.1. The first proof
Let $X$ be a $\operatorname{CAT}(\kappa)$ surface and let $x\in X$. Let $r>0$ be sufficiently small so that $B(x,r)$ is $\operatorname{CAT}(\kappa)$ and convex in $X$. We assume without loss of generality that $\kappa=1$. Use 2.5 to find a triangulation $\mathcal{S}$ of $B(x,r)$ with mesh less than $2D_{\kappa}$. Let $S_{1},\ldots,S_{n}$ be the collection of triangles in $\mathcal{S}$ that contain $x$. Then $S=\bigcup_{i=1}^{n}S_{i}$ is a closed topological disk such that $x\in S^{\circ}$. This follows from the convexity of each triangle and uniqueness of geodesics. We claim that $S^{\circ}$ is the neighborhood of $x$ as required by the definition of surface of bounded curvature.
Let $\mathcal{T}=\{T_{j}\}_{j=1}^{m}$ be an arbitrary collection of simple non-overlapping triangles in $S^{\circ}$. Our goal is to find a uniform upper bound on $\delta(\mathcal{T})$. We apply 4.4 to find for each $T_{j}\in\mathcal{T}$ a triangulation $\mathcal{T}_{j}$ of $T_{j}$ that is a vertex-edge refinement of the collection of polygons $S_{i}\cap T_{j}$, $i=1,\ldots,n$, intersecting the interior of $T_{j}$ with respect to $T_{j}$. Note that 2.3 implies that each set $S_{i}\cap T_{j}$ is in fact a polygon provided that its interior is non-empty. Observe from 4.5 that $\delta(T_{j})\leq\delta(\mathcal{T}_{j})$.
Next, for each triangle $S_{i}$, let $\mathcal{S}_{i}$ be the set of triangles in $\bigcup_{j}\mathcal{T}_{j}$ contained in $S_{i}$. We apply 4.4 a second time to $\mathcal{S}_{i}$ to find a vertex-edge refinement of $\mathcal{S}_{i}$ with respect to $S_{i}$. Denote this by $\mathcal{S}_{i}^{\prime}$. Suppose that $T\in\mathcal{T}_{j}$ and contained in $S_{i}$. Let $\mathcal{S}_{i}^{\prime}(T)$ denote the subset of $\mathcal{S}_{i}^{\prime}$ contained in $T$. Then, according to 4.4, $\mathcal{S}_{i}^{\prime}(T)$ is a vertex-edge triangulation of $T$ and hence $\delta(T)\leq\delta(\mathcal{S}_{i}^{\prime}(T))$ by 4.5. This gives
$\sum_{j=1}^{m}\delta(T_{j})$ $\leq\sum_{j=1}^{m}\sum_{T\in\mathcal{T}_{j}}\delta(T)\leq\sum_{i=1}^{n}\sum_{T\in\mathcal{S}_{i}}\sum_{T^{\prime}\in\mathcal{S}_{i}^{\prime}(T)}\delta(T^{\prime})\leq\sum_{i=1}^{n}\sum_{T\in\mathcal{S}_{i}}\sum_{T^{\prime}\in\mathcal{S}_{i}^{\prime}(T)}|\bar{T}_{1}^{\prime}|.$
Here, $\bar{T}_{1}^{\prime}$ is the model triangle in $M_{1}^{2}$ corresponding to $T^{\prime}$. Similarly, for each triangle $S_{i}$ we let $\bar{S}_{i}$ denote the model triangle of $S_{i}$ in $M_{1}^{2}$. Since $\mathcal{S}_{i}^{\prime}$ is a vertex-edge refinement with respect to $S_{i}$, 4.5 gives
$\sum_{i=1}^{n}\sum_{T\in\mathcal{S}_{i}}\sum_{T^{\prime}\in\mathcal{S}_{i}^{\prime}(T)}|\bar{T}_{1}^{\prime}|\leq\sum_{i=1}^{n}\sum_{T^{\prime}\in\mathcal{S}_{i}^{\prime}}|\bar{T}_{1}^{\prime}|\leq\sum_{i=1}^{n}|\bar{S}_{i}|.$
Combining these inequalities gives an upper bound for $\delta(\mathcal{T})=\sum_{j=1}^{m}\delta(T_{j})$ independent of the choice of $\mathcal{T}$. This completes the proof.
5.2. The second proof
Let $(X,d)$ be a $\operatorname{CAT}(\kappa)$ surface. Without loss of generality, we assume that $\kappa=1$. Given a point $x\in X$, let $U$ be a neighborhood of $x$ with compact closure. As shown in Section 3.4, for such a neighborhood $|U|$ is finite; note that this is the only step in this proof that relies on 2.5 concerning triangulations. We may choose $U$ to be sufficiently small so that $\delta(T)\leq\pi/20$ for any triangle $T\subset U$, 3.1 holds for $\theta=\pi/20$, and the model triangles $\bar{T}_{0}$ and $\bar{T}_{1}$ are $2$-bi-Lipschitz equivalent for any triangle $T\subset U$.
To verify that $X$ has bounded curvature, it is enough to show that there is a constant $C\geq 1$ such that
(5.1) $\delta(T)\leq C|T|$
for all simple triangles $T\subset U$. Let $T$ be such a triangle. If $\delta(T)\leq 0$, then the claim holds trivially. Thus we may assume that $\delta(T)>0$. Ideally, we would be able to obtain (5.1) as a consequence of 3.1. The obstacle is that $T$ may have a small angle. The strategy of our proof is to subdivide $T$ into two parts: a collection of triangles to which 3.1 applies and at most two triangles of small diameter.
Denote the angles of $T$ by $\alpha,\beta,\gamma$, arranged so that $\alpha\geq\beta\geq\gamma$. Let $p,q,r$ denote the vertices at $\alpha,\beta,\gamma$, respectively, and $A,B,C$ the edges opposite $\alpha,\beta,\gamma$, respectively. Observe in particular that $\alpha\geq\pi/3$. In the case that $\gamma\geq\pi/20$, we apply 3.1 (3) to conclude that $|T|\geq\delta(T)/(8L^{2})$ for some fixed constant $L$. This verifies (5.1) in the special case that $\gamma\geq\pi/20$.
[…]
Figure 5.1. The triangle $T$ in the second proof, following reductions
Thus we assume in the following that $\gamma\leq\pi/20$. As a consequence, observe that $\alpha\in[9\pi/20,\pi)$ (the possibility $\alpha=\pi$ violates the uniqueness of geodesics for $\operatorname{CAT}(\kappa)$ surfaces). In the case that $\alpha\geq 14\pi/20$, we perform an initial subdivision of $T$. Let $E$ be an arc that bisects the angle $\alpha$ and intersects $A$ at a point $e$. This forms two subtriangles $T_{1}=[pqe]$ and $T_{2}=[pre]$. Let $\alpha_{1}$ denote the angle at $p$ in the triangle $T_{1}$ and $\alpha_{2}$ the same for $T_{2}$. Then $\alpha_{1},\alpha_{2}\in[7\pi/20,\pi/2)$. Note that $\beta\leq 6\pi/20$. We see that for $T_{1}$, one of the three possibilities holds: the excess $\delta(T_{1})$ is negative; each angle of $T_{1}$ is at least $\pi/20$; or, the angle $\beta$ is at most $\pi/20$. The first two cases are handled by the previous observations. In turn, for the triangle $T_{2}$, we already have that $\gamma\leq\pi/20$. Since $\delta(T)\leq\delta(T_{1})+\delta(T_{2})$, it is enough to prove (5.1) for the triangles $T_{1}$ and $T_{2}$ separately.
In this way, we reduce to the case that $T$ has largest angle $\alpha\in[9\pi/20,14\pi/20]$ and smallest angle $\gamma\leq\pi/20$. Observe as a consequence that $\beta\in[5\pi/20,12\pi/20]$. See Figure 5.1.
We apply the following procedure to split off good triangles from $T$. Assume we have a triangle $T=T_{0}^{\prime}$ such that $\alpha,\beta\in[4\pi/20,17\pi/20]$. By relabeling vertices if needed, we may assume that $\alpha\geq\beta\geq\gamma$, and in particular (given that $\delta(T)>0$) that $\alpha\geq 9\pi/20$. Let $C_{1}$ be the arc bisecting the angle $\alpha$, which intersects $A$ at a point $q_{2}$. This splits $T$ into two triangles $T_{1}=[pq_{2}q]$ and $T_{1}^{\prime}=[pq_{2}r]$. Let $\beta_{1}$ denote the angle at $q_{2}$ in $T_{1}$ and $\beta_{1}^{\prime}$ denote the angle at $q_{2}$ in $T_{1}^{\prime}$. Define $\alpha_{1},\alpha_{1}^{\prime}$ similarly. Note that $\alpha_{1}^{\prime}\geq 4\pi/20$, which implies that $\beta_{1}^{\prime}\leq 17\pi/20$. This in turn implies that $\beta_{1}\geq 3\pi/20$. Thus all the angles of $T_{1}$ are at least $203\pi/$ 20, and we can apply 3.1 to $T_{1}$. On the other hand, the triangle $T_{1}^{\prime}$ satisfies $\alpha_{1}^{\prime},\beta_{1}^{\prime}\in[4\pi/20,17\pi/20]$, the same angle conditions as $T$ itself. Let $p_{1}$ denote the vertex of $T_{1}$ lying on the edge $B$ and $q_{1}$ denote the vertex on the edge $A$. Thus we may repeat the procedure to obtain triangles $T_{2},T_{3},T_{4},T_{5},\ldots$, vertices $p_{2},q_{2},p_{3},q_{3},\ldots$, where each $p_{n}$ lies on $B$ and each $q_{n}$ lies on $A$. By 3.1, the lengths of each of the edges of $T_{n}$ are comparable.
We need to show that we can terminate this process after finitely many steps. More precisely, we claim that for all $\varepsilon>0$ one can find $n\in\mathbb{N}$ such that $\delta(T_{n})<\varepsilon$, which follows if we can show that $\operatorname{diam}(T_{n}^{\prime})\to 0$ as $n\to\infty$. First, by the comparability of the edge lengths of $T_{n}$, we can rule out the case that $p_{n}\to r$ but $q_{n}$ does not, or vice versa. Next, the same comparability rules out the case that both $p_{n}$ and $q_{n}$ converge to a limit point other than $r$. We conclude that both $p_{n}$ and $q_{n}$ converge to $r$. This implies that $\operatorname{diam}T_{n}^{\prime}\to 0$.
5.3. The third proof
Let $X$ be a locally $\operatorname{CAT}(\kappa)$ surface. Let $\mathcal{U}$ be a cover of $X$ by convex open sets of diameter less than $2D_{\kappa}$ on which the restriction of the metric is $\operatorname{CAT}(\kappa)$. Let $\varepsilon_{n}$ be a sequence of decreasing reals each less than $2D_{\kappa}$ and limiting to 00. Apply 2.5 to obtain a sequence of triangulations $\mathcal{T}_{n}$ of $X$ corresponding to the parameters $\varepsilon_{n}$ and the collection $\mathcal{U}$; in particular, each $\mathcal{T}_{n}$ has mesh less than $2D_{\kappa}$. By 4.4, we can choose the $\mathcal{T}_{n}$ so that each triangle $S\in\mathcal{T}_{n}$ for all $n\geq 2$ is contained in some triangle of $\mathcal{T}_{1}$, and so that for each $T\in\mathcal{T}_{1}$ the subset of $\mathcal{T}_{n}$ of triangles intersecting the interior of $T$ form a vertex-edge triangulation of $T$.
Let $\mathcal{E}_{n}$ be the edge graph of $\mathcal{T}_{n}$. According to 2.5, the induced length metric $\widetilde{d}_{n}$ on $\mathcal{E}_{n}$ satisfies
$d(x,y)\leq\widetilde{d}_{n}(x,y)\leq d(x,y)+\varepsilon_{n}$
for all $x,y\in\mathcal{E}_{n}$. Define the surface $X_{n}$ by, for each $T\in\mathcal{T}_{n}$, gluing in a model triangle of constant curvature $\kappa$ having the same edge lengths into the edge graph $\mathcal{E}_{n}$. Identify $\mathcal{E}_{n}$ with a subset of $X_{n}$ in the natural way, and $X_{n}$ with $X$ by choosing some bijection that is the identity map on the edge graph $\mathcal{E}_{n}$. Having made these identifications, denote the new metric on $X$ by $\bar{d}_{n}$.
It follows from the $\operatorname{CAT}(\kappa)$ condition that $d(x,y)\leq\bar{d}_{n}(x,y)$ for all $x,y\in\partial T$ for some triangle $T\in\mathcal{T}_{n}$. Given two points $x,y\in X$, let $\bar{\Gamma}_{n}$ be a curve from $x$ to $y$ whose $\bar{d}_{n}$-length satisfies $\ell_{\bar{d}_{n}}(\Gamma_{n})\leq\bar{d}_{n}(x,y)+\varepsilon_{n}$. We assume that the restriction of $\bar{\Gamma}_{n}$ to each triangle $T\in\mathcal{T}_{n}$ is a $\bar{d}_{n}$-geodesic. Define the curve $\Gamma$ by replacing the subcurve of $\bar{\Gamma}_{n}$ intersecting a given triangle $T$ with the $d$-geodesic connecting the same endpoints. Then $\Gamma$ satisfies $\ell_{d}(\Gamma)\leq\ell_{\bar{d}_{n}}(\bar{\Gamma}_{n})+2\varepsilon_{n}$. Combining these inequalities gives
$d(x,y)\leq\ell_{d}(\Gamma)\leq\ell_{\bar{d}_{n}}(\bar{\Gamma}_{n})+2\varepsilon_{n}\leq\bar{d}_{n}(x,y)+3\varepsilon_{n}.$
Next, for any points $x,y\in X$, we can find points $x^{\prime},y^{\prime}\in\mathcal{E}_{n}$ such that $\bar{d}_{n}(x,x^{\prime})\leq\varepsilon_{n}$ and $\bar{d}_{n}(y,y^{\prime})\leq\varepsilon_{n}$. So for all $x,y\in X$, we have
$\bar{d}_{n}(x,y)\leq\bar{d}_{n}(x^{\prime},y^{\prime})+2\varepsilon_{n}\leq\widetilde{d}_{n}(x^{\prime},y^{\prime})+2\varepsilon_{n}\leq d(x^{\prime},y^{\prime})+3\varepsilon_{n}\leq d(x,y)+5\varepsilon_{n}.$
We conclude that the spaces $X_{n}$ converge uniformly to $X$.
Finally, we claim that the surfaces $X_{n}$ have locally uniformly bounded curvature. Let $x\in X$. Let $V$ be the interior of the union of all triangles in $\mathcal{T}_{1}$ containing $x$, observing that $x\in V$ and that its closure $\overline{V}$ is a closed polygon containing $x$. For all $n\in\mathbb{N}$, let $V_{n}$ denote the set $V$ considered with the metric $\bar{d}_{n}$. Note that each vertex of $X_{n}$ has total angle at least $2\pi$ and hence negative curvature. So all the positive curvature on $X_{n}$ comes from the interior of the faces. Since $\mathcal{T}_{n}$ is a vertex-edge triangulation, we have the uniform bound $|V_{n}|\leq|V_{1}|$. We conclude the positive part of the curvature on $V_{n}$ is at most $\kappa|V_{1}|$. In summary, $X$ is the uniform limit of surfaces $X_{n}$ of locally uniformly bounded curvature, and hence $X$ itself has bounded curvature.
6. Excess-curvature inequality
In this section, we give a proof of 1.2 relating the excess and total curvature of an arbitrary $\operatorname{CAT}(\kappa)$ triangle. First we verify 1.3, that the Hausdorff area of a surface $X$ of bounded curvature without cusp points coincides with the Alexandrov area. This is an application of Reshetnyak’s theorem that the metric tangent at each point is a cone over a circle. Recall that, for a Borel set $A\subset X$, we use $|A|$ to denote the Hausdorff area (i.e., $2$-measure) and $\text{Area}(A)$ to denote Alexandrov area.
Proof of 1.3.
We must show that $|A|=\text{Area}(A)$ for an arbitrary Borel set $A$. Since both Hausdorff area and Alexandrov area are determined by their values on polygons, without loss of generality we may assume that $A$ is a polygon. Let $A\subset X$ be a polygon. We can consider $A$ as its own metric space with the induced length metric; this does not affect either notion of area. Let $\varepsilon>0$ be given. For each $x\in X$, let $B_{x}$ be an open metric ball at $x$ with the property of being $(1+\varepsilon)$-bi-Lipschitz equivalent to a ball in a Euclidean cone over a circle, and let $\mathcal{U}=\{B_{x}:x\in A\}$. Let $\mathcal{T}$ be a triangulation of $A$ subordinate to the cover $\mathcal{U}$ and having mesh $\varepsilon$. Define a polyhedral surface $A_{\varepsilon}$ by replacing each triangle $T\in\mathcal{T}$ with the corresponding model triangle $\bar{T}_{0}$. It is clear that $|A_{\varepsilon}|=\text{Area}(A_{\varepsilon})$. Moreover, we have that each triangle $T\in\mathcal{T}$ is $(1+\varepsilon)$-bi-Lipschitz equivalent to the triangle $\bar{T}_{0}$, which implies that
$(1+\varepsilon)^{-2}|A|\leq|A_{\varepsilon}|\leq(1+\varepsilon)^{2}|A|.$
Letting $\varepsilon\to 0$, we have that $\text{Area}(A_{\varepsilon})$ converges to $\text{Area}(A)$; see Theorem VIII.2 in [1]. From this we conclude that $|A|=\text{Area}(A)$, which completes the proof. ∎
We now proceed with the proof of 1.2.
Proof of 1.2.
Let $T$ be a triangle with angles $\alpha,\beta,\gamma$. By subdividing if necessary, we may assume by 3.5 that $T$ is convex. In particular, we can now consider $T$ as its own metric space. Let $T_{n}$ be a sequence of model polyhedral approximations converging to $T$ as constructed in the third proof of 1.1; such a sequence can also be found by invoking the general theory of surfaces of bounded curvature. Note that $T_{n}$ is also a triangle, since the edges of $T$ are still geodesics in the polyhedral approximation. Let $\alpha_{n},\beta_{n},\gamma_{n}$ denote the angles in $T_{n}$ corresponding to the same edges in $T$ as $\alpha,\beta,\gamma$, respectively. The approximations $T_{n}$ come with a Lipschitz map $\varphi_{n}\colon T_{n}\to T$. This implies that $\alpha\leq\alpha_{n}$, $\beta\leq\beta_{n}$, $\gamma\leq\gamma_{n}$, and thus that $\delta(T)\leq\delta(T_{n})$.
The rotation of the boundary $\partial T_{n}$ is
$\tau(T_{n})=(\pi-\alpha_{n})+(\pi-\beta_{n})+(\pi-\gamma_{n})+\tau_{0}(T_{n})=2\pi-\delta(T_{n})+\tau_{0}(T_{n}),$
where $\tau_{0}(T_{n})$ is the contribution to the rotation of $\partial T_{n}$ from the interior points of the edges. The fact that each edge is a geodesic implies that $\tau_{0}(T_{n})\leq 0$. In particular, $\delta(T_{n})\leq 2\pi-\tau(T_{n})$. The Gauss–Bonnet theorem for surfaces of bounded curvature states that $\tau(T_{n})+\omega(T_{n})=2\pi$, where $\omega(T_{n})$ is the interior curvature of $T_{n}$. This relation gives
$\delta(T_{n})\leq 2\pi-\tau(T_{n})=\omega(T_{n}).$
Now, the positive part $\omega_{+}(T_{n})$ satisfies $\omega_{+}(T_{n})=\kappa|T_{n}|$. Since $\omega(T_{n})\leq\omega_{+}(T_{n})$ , we conclude that
$\delta(T)\leq\delta(T_{n})\leq\kappa|T_{n}|.$
Finally, the sequence $|T_{n}|$ converges to $|T|$ as $n\to\infty$. This completes the proof. ∎
7. Approximation by smooth Riemannian surfaces
In this section, we prove that every $\operatorname{CAT}(\kappa)$ surface, $\kappa\in\mathbb{R}$, is the uniform limit of smooth Riemannian surfaces with Gaussian curvature bounded above by $\kappa$ and having locally uniformly bounded area. In Section 5.3, we showed that such a surface is the uniform limit of $\operatorname{CAT}(\kappa)$ model polyhedral surfaces. What remains is to show that one can smoothen out the vertices while retaining the $\operatorname{CAT}(\kappa)$ condition.
Let $X$ be a polyhedron with faces of constant curvature $\kappa$. The metric in a neighborhood of a vertex, in polar coordinates, is defined by the conformal length element $ds^{2}=\lambda(r)(dr^{2}+r^{2}d\theta^{2})$, where
(7.1) $\lambda(r)=\frac{4\alpha^{2}r^{2(\alpha-1)}}{(1+\kappa\cdot r^{2\alpha})^{2}}$
and $\alpha\geq 1$ is the total angle at the vertex divided by $2\pi$. If $\alpha=1$, then we just have the usual constant-curvature metric. Observe that $\lim_{r\to 0}\lambda(r)=0$ if $\alpha>1$. The Gaussian curvature of this metric outside of the vertex is given by the formula
$K(r)=\frac{-1}{2\lambda(r)}\left(\frac{\partial^{2}}{\partial r^{2}}\log(\lambda(r))+\frac{1}{r}\frac{\partial}{\partial r}\log(\lambda(r))\right).$
We have the following main lemma.
Lemma 7.1.
Consider the Euclidean ball $B_{\operatorname{Euc}}(\mathbf{0},R)\subset\mathbb{R}^{2}$ for some $R>0$, equipped with the metric $ds^{2}=\lambda(r)(dr^{2}+r^{2}d\theta^{2})$, with $\lambda$ given by (7.1), with $\alpha>1$ and $\kappa\geq 0$. For all $\delta\in(0,R)$ sufficiently small, there is a smooth Riemannian metric $ds_{\delta}^{2}=\lambda_{\delta}(r)(dr^{2}+r^{2}d\theta^{2})$ that agrees with $ds^{2}$ on $B_{\operatorname{Euc}}(\mathbf{0},R)$ outside the Euclidean ball $B_{\operatorname{Euc}}(\mathbf{0},\delta)$ and has Gaussian curvature at most $\kappa$. Moreover, $\lambda_{\delta}$ is an increasing function on $B_{\operatorname{Euc}}(\mathbf{0},\delta)$.
Proof.
We give two approaches to proving this lemma: the first based on interpolating with a flat surface and the second based on interpolating with a hyperbolic surface. The first approach is simpler but only covers the case where $\kappa\geq 0$.
1. Interpolating with a flat metric. We assume here that $\kappa\geq 0$. We define the smoothened metric $\lambda_{\delta}$ as follows. First, we define for each $0<a<b$ the smooth cutoff function $\varphi_{a,b}:[0,\infty)\to[0,1]$ by
$\varphi_{a,b}(r)=\left\{\begin{array}[]{cl}0&\text{ if }r\leq a\\ \frac{e^{-1/(r-a)}}{e^{-1/(b-r)}+e^{-1/(r-a)}}&\text{ if }a\leq r\leq b\\ 1&\text{ if }r\geq b\end{array}\right..$
Write $\varphi_{b}=\varphi_{b/2,b}$. Next, let
$g_{\delta}(r)=\left(\frac{\partial}{\partial r}\log(\lambda(r))\right)\varphi_{\delta}(r)$
and
[…]
We claim that this choice of $\lambda_{\delta}$ gives the required properties. It is easy to see that $\lambda_{\delta}(r)=\lambda(r)$ whenever $r\geq\delta$. Moreover, $g_{\delta}(r)=0$ if $r\leq\delta/2$, from which it follows that $\lambda_{\delta}$ is constant on the interval $[0,\delta]$. From this, we see that $ds_{\delta}$ is locally isometric to the plane on $B(\mathbf{0},\delta)$. We conclude that $ds_{\delta}$ is a smooth Riemannian metric. Observe now that
(7.2) $\frac{\partial}{\partial r}\log(\lambda(r))=\frac{-2+2\alpha-2\kappa\cdot r^{2\alpha}(1+\alpha)}{r(1+\kappa\cdot r^{2\alpha})},$
which is positive provided that $\alpha>0$ and $r$ is sufficiently small depending on $\alpha$. Take $\delta>0$ small enough so that the quantity in (7.2) is positive for all $r\in(0,\delta)$. It follows that $g_{\delta}(r)$ is non-negative for all $r\in(0,\delta)$ and thus that $\lambda_{\delta}(r)$ is increasing on this interval. We also choose $\delta$ small enough so that $\lambda(\delta)\leq 1$.
The curvature of the metric $ds_{\delta}^{2}$ is
$K_{\delta}(r)$ $=\frac{-1}{2\lambda_{\delta}(r)}\left(g_{\delta}^{\prime}(r)+\frac{g_{\delta}(r)}{r}\right)$
$=\frac{-1}{2\lambda_{\delta}(r)}\left(\left(\frac{\partial}{\partial r}\log(\lambda(r))\right)\varphi_{\delta}^{\prime}(r)+\left(\frac{\partial^{2}}{\partial r^{2}}\log(\lambda(r))\right)\varphi_{\delta}(r)+\frac{\left(\frac{\partial}{\partial r}\log(\lambda(r))\right)\varphi_{\delta}(r)}{r}\right).$
To complete the proof, we must show that $K_{\delta}(r)\leq\kappa$. This is equivalent to the inequality
(7.3) $-g_{\delta}^{\prime}(r)-\frac{g_{\delta}(r)}{r}\leq 2\kappa\cdot\lambda_{\delta}(r),$
which we can write as
[…]
To handle the first term on the left-hand side of this equation, observe that $\left(\frac{\partial}{\partial r}\log(\lambda(r))\right)\varphi_{\delta}^{\prime}(r)\geq 0$, again using the property that $\frac{\partial}{\partial r}\log(\lambda(r))>0$ for all $r\in(0,\delta)$. The other two terms can be written as
$\left(-\frac{\partial^{2}}{\partial r^{2}}\log(\lambda(r))-\frac{\left(\frac{\partial}{\partial r}\log(\lambda(r))\right)}{r}\right)\varphi_{\delta}(r)=2\kappa\cdot\lambda(r)\varphi_{\delta}(r).$
Thus, to verify (7.3), it is enough to show that $2\kappa\cdot\lambda(r)\varphi_{\delta}(r)\leq 2\kappa\cdot\lambda_{\delta}(r)$ for all $r\in(\delta/2,\delta)$. This is trivially satisfied if $\kappa=0$. If $\kappa>0$, this can be written as
[…]
First, since $\lambda(r)\leq\lambda(\delta)\leq 1$ and $\varphi_{\delta}(r)\leq 1$ for all $r\in(0,\delta)$, we have
$\lambda(r)\varphi_{\delta}(r)\leq\lambda(r)\leq\lambda(r)^{\varphi_{\delta}(r)}=\exp\left((\log\lambda(r))\varphi_{\delta}(r)\right).$
Then, using the properties that $r<\delta$ and that $\varphi_{\delta}$ is increasing and bounded by $1$, we obtain
[…]
This concludes our verification that $K_{\delta}(r)\leq\kappa$.
2. Interpolating with the hyperbolic metric. The following proof handles the general case. It is similar to the previous one, so we focus on the differences. By rescaling, we may assume that $\kappa\geq-1$. Let $\bar{\lambda}(r)=4/(1-r^{2})^{2}$ be the conformal weight for the hyperbolic metric, which is defined for all $r\in(0,1)$. We make two immediate observations. First, since $\alpha>1$ and $\kappa\geq-1$, we have $\bar{\lambda}(r)\geq\lambda(r)$ for all $r\in(0,1)$. Second, the inequality
(7.5) $\frac{\partial}{\partial r}\log(\lambda(r))\geq\frac{\partial}{\partial r}\log(\bar{\lambda}(r))$
holds for all $r>0$ sufficiently small. Assume then that $\delta>0$ is small enough that (7.5) holds for all $r\leq\delta$. We also necessarily assume that $\delta<1$.
Our main modification is to redefine $g_{\delta}$ above using the formula
$g_{\delta}(r)=\left(\frac{\partial}{\partial r}\log(\lambda(r))\right)\varphi_{\delta}(r)+\left(\frac{\partial}{\partial r}\log(\bar{\lambda}(r))\right)(1-\varphi_{\delta}(r)).$
[…]
Next, to verify (7.3), there are two details to check. First, we observe from (7.5) that
$\left(\frac{\partial}{\partial r}\log(\lambda(r))-\frac{\partial}{\partial r}\log(\bar{\lambda}(r))\right)\varphi_{\delta}^{\prime}(r)\geq 0$
for all $r\in(0,\delta)$. Then, instead of (7.4) we must check that
$\kappa\cdot\lambda(r)\varphi_{\delta}(r)-\bar{\lambda}(r)(1-\varphi_{\delta}(r))\leq\kappa\cdot\lambda_{\delta}(r)$
for all $r\in(0,\delta)$. This is trivial if $\kappa=0$. If $\kappa>0$, it suffices to show that
$\lambda(r)\varphi_{\delta}(r)\leq\lambda_{\delta}(r),$
which can be done similarly to the first part of the proof. In the case where $\kappa<0$, it suffices to show that
$\lambda(r)\varphi_{\delta}(r)+\bar{\lambda}(r)(1-\varphi_{\delta}(r))\geq\lambda(\delta)\exp\left(\int_{\delta}^{r}g_{\delta}(t)\right).$
Note that, since $\kappa<0$, this inequality is reversed relative to (7.4). We first use Young’s inequality to obtain
$\lambda(r)\varphi_{\delta}(r)+\bar{\lambda}(r)(1-\varphi_{\delta}(r))\geq\exp\left((\log\lambda(r))\varphi_{\delta}(r)+(\log\bar{\lambda}(r))(1-\varphi_{\delta}(r))\right).$
The right-hand side can be written as
[…]
Since $\bar{\lambda}(r)\geq\lambda(r)$ and $\varphi_{\delta}$ is increasing, this quantity is greater than
[…]
This concludes the proof. ∎
We now proceed with the proof of 1.4.
Proof of 1.4.
Let $X$ be a $\operatorname{CAT}(\kappa)$ surface, where $\kappa\geq 0$. Let $(\varepsilon_{n})$ be a sequence of positive reals limiting to 00, and $X_{n}$ the approximating model polyhedral surface given in Section 5.3 for the value $\varepsilon_{n}$. In particular, there is a homeomorphic $\varepsilon_{n}$-isometry $\psi_{n}\colon X_{n}\to X$ with the property that $|\psi_{n}^{-1}(A)|$ converges to $|A|$ for all sets $A\subset X$. For each vertex $v_{i}$ of $X_{n}$, $i\in\mathbb{N}$, let $r_{i}>0$ be such that $B(v_{i},2r_{i})$ does not contain any other vertex of $X_{n}$. Note that this implies that the balls $B(v_{i},r_{i})$ are disjoint. We also require that $\sum_{i=1}^{\infty}r_{i}<\varepsilon_{n}$. The ball $B(v_{i},r_{i})$ is locally isometric to $B_{\operatorname{Euc}}(\mathbf{0},R_{i})$ for some $R_{i}>0$, equipped with the metric given in 7.1. Choose $\delta_{i}>0$ sufficiently small so that $B_{\operatorname{Euc}}(\mathbf{0},\delta_{i})$ has diameter and area at most $\varepsilon_{n}2^{-i}$, which can be done since $\lambda_{\delta}$ is increasing for $r\in(0,\delta)$ and $\lambda(\delta)\to 0$ as $\delta\to 0$. Let $Y_{n}$ be the smooth Riemannian surface obtained by applying 7.1 to each neighborhood $B_{\operatorname{Euc}}(\mathbf{0},\delta_{i})$. Let $V_{i}$ denote the neighborhood in $X_{n}$ on which the metric is altered.
Define a map $\varphi_{n}$ from $X_{n}$ to $Y_{n}$ by mapping $V_{i}$ to the corresponding neighborhood of $Y_{n}$ in a continuous way fixing $\partial V_{i}$ and as the identity map otherwise. By construction, $\varphi_{n}$ is a $(2\varepsilon_{n})$-isometry. Moreover, for any set $A\subset X_{n}$, we have $|\varphi_{n}(A)|\leq|A|+\varepsilon_{n}$. Then $\psi_{n}\circ\varphi_{n}^{-1}$ is a $(3\varepsilon)$-isometry from $Y_{n}$ to $X$ with the property that $|\varphi_{n}\circ\psi_{n}^{-1}(A)|$ converges to $|A|$ for all sets $A\subset X$. ∎
Remark 7.2.
The ideas of 7.1 also apply to smoothing vertices of model polyhedral surfaces with curvature bounded below. Assume that $X$ is a model polyhedral surface with curvature bounded below for some $\kappa\leq 0$. Then the metric is given in a neighborhood of a vertex by (7.1), where $\alpha<1$. Thus, for sufficiently small $r$, one has $\frac{\partial}{\partial r}\log(\lambda(r))>0$ by (7.2). Since $\kappa\leq 0$, the chain of inequalities at the end of the first part of the proof of 7.1 establishes (7.3) with $\leq$” replaced by $\geq$”. We conclude that the Riemannian approximation in 7.1 has curvature bounded below by $\kappa$.
The general case can be handled by interpolating with the spherical metric, by modifying the above argument for the hyperbolic metric. We leave the details to the interested reader.